Carlos Company had the following stock outstanding and Retained Earnings at December 31, 2018:

On December 31, 2018, the board of directors is considering the distribution of a cash dividend to the common and preferred stockholders. No dividends were declared during 2016 or 2017. Three independent cases are assumed:

Case A: The preferred stock is noncumulative; the total amount of 2018 dividends would be $22,000.
Case B: The preferred stock is cumulative; the total amount of 2018 dividends would be $22,000. Dividends were not in arrears prior to 2016.
Case C: Same as Case B, except the amount is $65,000.


Required:

1-a. Compute the amount of 2018 dividends, in total, that would be payable to each class of stockholders if dividends were declared as described in each case.

1-b. Compute the amount of 2018 dividends per share payable to each class of stockholders for each case.
Answer #1

Preferred dividend = 188000*8% = 15040

1a) Dividend distribution

Case A Case B Case C
Preferred 15040 22000 15040*3 = 45120
Common 6960 0 65000-45120 = 19880

Preferred dividend = 188000*8% = 15040

1b) Dividend distribution

Case A Case B Case C
Preferred 0.80 1.70 2.40
Common 0.01 0 0.04